# Immunity to Increasing Condition Numbers of Linear Superiorization versus Linear Programming

### v1

@@ 4 $x^{-1}\leftarrow x^{0}+\mathbbm{1}$;
5 while $\max_{i}(\langle a^{i},x^{k}\rangle-b_{i})\geq\epsilon$ or $\frac{\|x^{k}-x^{k-1}\|}{\|x^{k-1}\|}\geq\epsilon$ do
@@ 11      $\tau\leftarrow\tau+1$;
12       if $\tau<\tau_{\text{reset}}$ then
@@ 22      $k\leftarrow k+1$;
3.3 The LP algorithms
We compare the superiorization method with the following LP algorithms:
@@ 4 Experimental Results of the Comparative Investigation
Our main finding is, that, based on our computational experiments, the superiorization methodology is quite robust with respect to increasing condition numbers. In particular, the runtime remains relatively small when compared to the scipy implementations of the simplex algorithm. We conjecture that this is due to the fact, that our superiorization implementation never considers the full problem at once, but performs individual projections onto the half-spaces instead. This comes at the cost of sacrificing feasibility during early iterations. Furthermore, we conjecture, that the bounded perturbation resilience of the basic algorithm may play a role in absorbing errors that occur during computation, which would normally be amplified by the condition number.
Depending on the rate at which the step-sizes of the perturbations converge to zero, this would mean that superiorization, in general, may be less affected by high condition numbers, than other current algorithms. A secondary finding is, that superiorization can produce better results than the scipy simplex implementations, when both are terminated at a certain time, that is before the usual termination criterion is met, cf. Figure 1(c). Furthermore, we see that superiorization reaches its termination criterion considerably faster than scipy in higher dimensions.
The trend is clear: While in problems with smaller dimensions the simplex algorithms are considerably faster, with increasing problem dimensions the scipy implementations take much longer to terminate. We also see a clear difference in the algorithms styles: Simplex aims to become feasible first and then starts to improve the objective function. Superiorization, on the other hand, reduces objective function values in its initial iterations because then the step-sizes of the perturbations are still large and only as iterations proceed the effect of feasibility-seeking becomes stronger.
Finally, it should be noted that scipy’s revised simplex performs better than scipy’s standard simplex method. This can be attributed to the internal $LU$ decomposition of the basic matrices, leading to better numerical stability as described in [4].
@@
(c) $400\times 500$. Had both implementations been terminated at an earlier time (say after 25 seconds), the superiorized solution would be both less violating the constraints and closer to optimality. As the problem dimensions increase this trend becomes more and more prominent
@@
Figure 1: Figures 1(a)-1(d) show the behavior of the different algorithms on problems with increasing dimension for fixed condition number $\kappa=1000$. On the left we see the maximum violation of the constraints $\max_{i}(\langle a^{i},x\rangle-b_{i})$ plotted against the runtime. On the right we have the corresponding objective function values $\langle c,x\rangle$ plotted against the runtime.
(a) $400\times 500$
(b) $800\times 1000$
(c) $2000\times 2500$
Figure 2: Figures 2(a)-2(c) demonstrate well the robustness with respect to increasing condition numbers. On the left we see the runtimes of each algorithm, plotted against the condition numbers, on the right we see the objective function value at termination, plotted against the condition numbers. Notice the severely suboptimal objective function value of scipy’s simplex for high condition numbers. This explains the low runtime as the algorithm realizes that these problems are hard and quickly “surrenders”. This trend continues in Figures 2(b) and 2(c) and the revised simplex too starts to reach its limits. Superiorization on the other hand proves to be quite stable in terms of its runtime with regards to increasing condition numbers.
We observe that in our experiments the interior point method and the Gurobi implementation outperformed the other algorithms by a lot. For the interior point method this matched our expectations and can be explained in the following way: The condition number of a matrix can be interpreted as a measure of how linearly dependent its rows or columns are. A well-conditioned matrix (i. e., $\kappa=1$) only has a single singular value and will be semi-orthogonal, whereas an ill-conditioned matrix ($\kappa$ “large”) will have “almost” linearly dependent entries. Consequently, the half-spaces $\{x\in\mathbb{R}^{n}\colon\langle a^{i},x\rangle\leq b_{i}\}$ will be almost parallel and the resulting polyhedron will consist of many facets and vertices. A basic simplex implementation, which moves from vertex to vertex, will consequently face a long runtime. The interior point method, on the other hand, is not dependent on the vertices. It will take its path through the interior of the polyhedron regardless of its boundary.
The Gurobi implementation is based on observations of the condition number. In that sense, our experiments played into one of Gurobis strongest suits. As one of the biggest commercial state of the art solvers we did not expect to outperform Gurobi with our implementation, but rather used it as a benchmark to compare with the other algorithms.
@@ 5 Conclusions
In this paper we experimentally discussed the superiorization method and constrained optimization on a set of examplary linear problems with varying condition numbers and problem sizes with the aim of investigating and comparing their immunity to increasing condition numbers. Our experimental results are promising for the observed problem sizes, but more work needs to be done for larger instances to verify that the trend that we observed continues.
The superiorization method and constrained optimization use the same input data which consists of a family of constraints obtained from the modeling process along with a user-chosen objective function. But the two approaches aim at different end-points of their iterative processes. The easy accessibility of the superiorization methodology allows for quick implementations with the advance knowledge that the aim is not to reach a constrained optimum. At the same time the SM may compute its solutions at a lower runtime (in the case of scipy’s simplex and revised simplex).
@@
Acknowledgements: The work of Y.C. is supported by U.S. National Institutes of Health Grant Number R01CA266467 and by the Cooperation Program in Cancer Research of the German Cancer Research Center (DKFZ) and Israel’s Ministry of Innovation, Science and Technology (MOST).

### v2

@@ 4 $x^{-1}\leftarrow x^{0}+\mathbbm{1}$;
5 while $\max_{i}(\langle a^{i},x^{k}\rangle-b_{i})\geq\varepsilon$ or $\frac{\|x^{k}-x^{k-1}\|}{\|x^{k-1}\|}\geq\varepsilon$ do
@@ 11 $\tau\leftarrow\tau+1$;
12 if $\tau<\tau_{\text{restart}}$ then
@@ 22 $k\leftarrow k+1$;
3.3 The LP solvers
We compared the actions of LinSup with those of the following LP solvers:
@@ 3.3 The LP solvers
3.4 Setting up the experiments
To compare the numerical performance of the different algorithms with each other in a fair manner we performed the following steps: First we ran the LinSup algorithm with an infeasibility threshold of $\varepsilon=10^{-8}$ as a stopping rule. We recorded the total run-time $T_{\text{sup}}$ as well as the run-times, objective function values and the maximal constraint violations at each iteration.
Then we turned toward the LP solvers: We imposed on each solver the same infeasibility threshold of $\varepsilon=10^{-8}$ as a stopping rule. Additionally we required the solver to stop, if it did not stop earlier, after the time $T_{\text{sup}}$ was exceeded. In this way, if the solver would not terminate according to the infeasibility threshold we stopped it after the time $T_{\text{sup}}$ was reached and then extracted the current iterate.
We then compared objective function values and maximal constraint violation of the LP solver with the corresponding values from the LinSup algorithm. Similarly, we also documented the run-times, objective function values and maximal constraint violations for each iteration of the LP solvers and plotted the run-times against objective function values and maximal constraint violations.
This experimental set up does not aim at reaching LP optimality because LinSup is not an LP solver. The aim is to reach a certain level of infeasibility (i.e., getting close to feasibility as we wish) and, at such a level, compare LinSup with the LP solvers with respect to the run-times needed to reach this level of infeasibility and by how they performed in reducing (not necessarily reaching optimality) the objective function value.
@@ 4 Experimental Results of the Comparative Investigation
We split the presentation of our results into two parts, first considering the effects of increasing problem sizes and second considering the effects of increasing condition numbers for fixed sizes.
4.1 Sensitivity to increasing sizes when the condition number is fixed
Our first finding is, that, based on our computational experiments, the superiorization methodology is quite robust with respect to increasing dimensions. In particular, the run-time remains relatively small when compared to the scipy implementations of the simplex algorithm.
Furthermore, we see that superiorization reaches its termination criterion considerably faster than the simplex implementations in higher dimensions. This shows that superiorization can produce better results than the simplex implementations, when both are terminated at a certain time, that is before the usual termination criterion is met, cf. Figure 1(c) and 2(a)2(c).
@@ 4.1 Sensitivity to increasing sizes when the condition number is fixed
(c) $400\times 500$. In this dimension the superiorization algorithm starts terminating before scipy’s simplex algorithms. We can see that at the time of termination, they have not even attained feasibility.
Figure 1: Dimensions $80\times 100$ - $400\times 500$. Figures 1(a)2(c) show the behavior of the different algorithms on problems with increasing dimension for fixed condition number $\kappa=1000$. On the left we see the maximum violation of the constraints $\max_{i}(\langle a^{i},x\rangle-b_{i})$ plotted against the run-time. On the right we have the corresponding objective function values $\langle c,x\rangle$ plotted against the run-time.
(a) $800\times 1000$
(b) $2000\times 2500$
(c) $4000\times 5000$. In this dimension we see that the superiorization algorithm terminates even before Gurobi’s simplex does.
Figure 2: Dimensions $800\times 1000$ - $4000\times 5000$.
The trend is clear: While in problems with smaller dimensions the simplex algorithms are considerably faster, with increasing problem dimensions the simplex implementations take much longer to terminate, up to the point where the superiorization method terminates quicker.
We also see a clear difference in the algorithms styles: Simplex aims to become feasible first and then starts to improve the objective function. Superiorization, on the other hand, reduces objective function values in its initial iterations because then the step-sizes of the perturbations are still large and only as iterations proceed the effect of feasibility-seeking becomes stronger.
Further, Gurobi outperformed both of the other simplex implementations, which is to be expected by a commercial product. Nevertheless, we can see that Gurobi, too, needed more and more time until termination with increasing dimension and eventually the superiorization algorithm terminated quicker.
4.2 Immunity to increased condition numbers
Our second finding is, that the superiorization methodology is able to easily handle large condition numbers, especially when compared to the scipy implementations. We conjecture that this is due to the fact, that our superiorization implementation never considers the full problem at once, but performs individual projections onto the half-spaces instead. This comes at the cost of sacrificing feasibility during early iterations.
Furthermore, we conjecture, that the bounded perturbation resilience of the basic algorithm may play a role in absorbing errors that occur during computation, which would normally be amplified by the condition number.
This would mean that superiorization, in general, may be less affected by high condition numbers, than other current algorithms.
We observe that in our experiments, in terms of runtime and objective function value, the interior point method seemed to outperformed the other algorithms. This can be explained in the following way:
The condition number of a matrix can be interpreted as a measure of how linearly dependent its rows or columns are. A well-conditioned matrix (i. e., $\kappa=1$) only has a single singular value and will be semi-orthogonal, whereas an ill-conditioned matrix ($\kappa$ “large”) will have “almost” linearly dependent entries. Consequently, the half-spaces $\{x\in\mathbb{R}^{n}\colon\langle a^{i},x\rangle\leq b_{i}\}$ will be almost parallel and the resulting polyhedron will consist of many facets and vertices.
A basic simplex implementation, which moves from vertex to vertex, will consequently face a long run-time. The interior point method, on the other hand, is not dependent on the vertices. It will take its path through the interior of the polyhedron regardless of its boundary.
(a) $80\times 100$
(b) $200\times 250$
(c) $400\times 500$
Figure 3: Figures 3(a)4(c) demonstrate well the robustness with respect to increasing condition numbers. On the left we see the averaged run-times of 55 problem instances for each algorithm, plotted against the condition numbers, on the right we see the averaged objective function values at termination, plotted against the condition numbers. Notice the severely suboptimal objective function value of scipy’s simplex for high condition numbers. This explains the low run-time as the algorithm realizes that these problems are hard and quickly “surrenders”.
(a) $800\times 1000$
(b) $2000\times 2500$
(c) $4000\times 5000$
Figure 4: The trend continues in Figures 4(a) and 4(c) and the revised simplex too starts to reach its limits. Superiorization on the other hand proves to be quite stable in terms of its run-time with regards to increasing condition numbers, at times surpassing the Gurobi implementation in both runtime and objective function value. While the total runtime does increase significantly with the condition number, it remains within reasonable timeframes, considering the increased difficulty of the problems that accompanies higher condition numbers
(a) $80\times 100$
(b) $200\times 250$
(c) $400\times 500$
@@ 4.2 Immunity to increased condition numbers
(e) $2000\times 2500$
(f) $4000\times 5000$
Figure 5: Average infeasibilities at termination of the problem instances. Notice how with increasing condition number the general trend is towards higher infeasibility, except for superiorization, which remains constant throughout. Like before, scipy.simplex and scipy.revised simplex turn out to be unreliable in moderate dimension already. Additionally, we can see that scipy.interior-point reaches infeasibilities in the order of magnitude $10^{0}-10^{1}$.
However, when observing the infeasibility at termination (Figures 5(a)-5(f)), we see that the interior point method has trouble attaining the desired level of feasibility. We can see values of infeasibility of magnitudes up to $10^{0}-10^{1}$, significantly increasing with the condition number. In other words, while the interior point method seems to be able to quickly find the correct objective function value, it has issues doing so in a feasible manner, making it unreliable for practical applications with high condition numbers.
On the other hand, superiorization consistently reaches precisely the required level of feasibility, regardless of the condition number. This becomes particularly evident when looking at the proximity-target-curves, i. e. the plot of the level of infeasibility against the objective function value for all iterates of the respective algorithms. Figures 6(a)-6(d) showcase the behaviour of the interior point method as well as of superiorization.
(a) $\kappa=1$
(b) $\kappa=10^{2}$
(c) $\kappa=10^{4}$
(d) $\kappa=10^{6}$
Figure 6: Proximity-Target-Curves of superiorization and of the interior-point method for multiple condition numbers. Notice once again the severe infeasibility (i.e. proximity) of some of the final iterates of the interior-point method.
Combining all of the above observations, superiorization clearly offers a valuable contribution to large-scale optimization problems of high condition numbers. For large-enough problem instances it is faster than scipy’s simplex methods and more reliable than scipy’s interior point method in terms of feasibility.
@@ 5 Conclusions
In this paper we experimentally discussed the superiorization method and constrained optimization on a set of exemplary linear problems with varying condition numbers and problem sizes with the aim of investigating and comparing their immunity to increasing condition numbers. It expands on the results in [12] and further investigates the strengths and weaknesses of superiorization and classical LP optimization algorithms. Our experimental results are promising for the observed problem sizes, and we are confident that the trend that we observed will continue for larger problem instances.
The superiorization method and constrained optimization use the same input data which consists of a family of constraints obtained from the modeling process along with a user-chosen objective function. But the two approaches aim at different end-points of their iterative processes. The easy accessibility of the superiorization methodology allows for quick implementations with the advance knowledge that the aim is not to reach a constrained optimum. At the same time the SM may compute its solutions at a lower run-time (in the case of simplex) and give better feasibility than the interior-point method.
@@
Statements and Declarations
Competing Interests: The authors have no competing interests to declare that are relevant to the content of this article.
All authors contributed to the study conception and design. All authors read and approved the final manuscript.
Data sets generated during the current study are available from the corresponding author on reasonable request.
Acknowledgments: The work of Y.C. is supported by U.S. National Institutes of Health Grant Number R01CA266467 and by the Cooperation Program in Cancer Research of the German Cancer Research Center (DKFZ) and Israel’s Ministry of Innovation, Science and Technology (MOST).